Pre lab questions
Draw two molecules that are examples of constituti
Answer #1

1. Here are two examples each of positional and skeletal isomers.

positional isomers skeletal isomers and and он он and Br and Br

2. The cyclohexane ring in boat and chair form is here:

chair boat

3. A plane of symmetry is an imaginary plane that divided the molecule in two halves. For example the molecule of mesotartaric acid has a plane of symmetry that divided the molecule in two halves and each of the half is the mirror image of the other.

COOH он он COOH, the plane of symmetry is shown by the dashed line.

Enantiomers do not have a plane of symmetry, if they had, they would be either achiral or meso compounds. See for yourself, the two enantiomers of lactic acid. Left one is (S)-lactic acid and the right one is (R)-lactic acid.

COOH HOOC OHHO

4. Optical activity is the rotation of plane of polarization of linear polarized light. The rotation occurs as the light passes through the sample of a chiral substance. Such a substance is called optically active. Remember, for a molecule to be optically active, it has to be asymmetric. If a molecule is symmetric, for example it contains a plane of symmetry, it will not be optically active. Optical activity, is measured by a polarimeter that uses a source of plane polarized light. The light passes through the sample and the plane of polarization rotates, the rotation is measured when the light passes through another polarizer.

the optical acitivty is denoted by \alpha and is given by:

\alpha = [\alpha]_{\lambda}^{T}\times l \times c, where [\alpha]_{\lambda}^{T} is the specific rotation at a fixed temperature and wavelength while l is the path length and c is the concentration of solution.

5. The priority order of groups is determined by the Cahn-Ingold-Prelog priority rules. The rules are:

the group or atom having the greater atomic number received higher priority. For example if oxygen and nitrogen both are attached to a carbon atom, oxygen atom will have the higher priority owing to its greater atomic number.

if there is a tie, the first point of difference rule holds true. For example this molecule

H2 CCI both the groups attached to carbon atom are CH2 but the group at the top has chlorine further ahead while the one at the bottom has bromine so the group at the bottom received a higher priority since bromine has the greater atomic number.

I hope it helps.